# Molecular relaxation by reverse diffusion with time step prediction

### v1

@@ Abstract
Molecular relaxation, finding the equilibrium state of a non-equilibrium structure, is an essential component of computational chemistry to understand reactivity. Classical force field methods often rely on insufficient local energy minimization, while neural network force field models require large labeled datasets encompassing both equilibrium and non-equilibrium structures. As a remedy, we propose MoreRed, molecular relaxation by reverse diffusion, a conceptually novel and purely statistical approach where non-equilibrium structures are treated as noisy instances of their corresponding equilibrium states. To enable the denoising of arbitrarily noisy inputs via a generative diffusion model, we further introduce a novel diffusion time step predictor. Notably, MoreRed learns a simpler pseudo potential energy surface instead of the complex physical potential energy surface. It is trained on a significantly smaller, and thus computationally cheaper, dataset consisting of solely unlabeled equilibrium structures, avoiding the computation of non-equilibrium structures altogether. We compare MoreRed to classical force fields, equivariant neural network force fields trained on a large dataset of equilibrium and non-equilibrium data, as well as a semi-empirical tight-binding model. To assess this quantitatively, we evaluate the root-mean-square deviation between the found equilibrium structures and the reference equilibrium structures as well as their DFT energies.
@@ 1 Introduction
Because iterative ab initio electronic structure calculations are computationally expensive and not feasible in high-throughput settings, methods for finding equilibrium structures need to be efficient [41, 42]. To address this limitation, numerous approaches have been developed that speed up the computation of forces but also suffer from a loss in accuracy compared to ab initio methods. These include classical force field (FF) methods like MMFF94 [43], the universal force field [44], or CHARMM [45], on the one hand, and semiempirical methods such as GFN2-xTB [46], PM6-7 [47, 48], or OM2 [49] on the other hand. Furthermore, machine learning FF (MLFF) models [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65] have emerged as promising alternatives to physical models. When trained on a sufficient amount of ab initio calculations, MLFF models, using kernel methods such as sGDML [66, 67, 68] or neural networks such as SchNet [69, 70], learn the physical PES and very efficiently predict the forces. Moreover, they have been shown to produce promising results in relaxation tasks [71, 72], while reducing computational cost by several magnitudes. Although MLFF models significantly accelerate gradient computations for structural relaxation, the training dataset must cover a wide range of the chemical space including equilibrium and non-equilibrium structures with accurately computed physical labels, introducing a large computational cost for generating a training dataset.
@@
An emerging machine learning-based approach to exploring chemical space is training generative models on a dataset of equilibrium structures in order to learn to generate new molecular structures. For instance, diffusion models have been used recently in molecule generation [73, 74, 75, 76, 77], conformer search [78] and molecular graph generation [79, 80]. They generate samples via iterative denoising, starting from a simple prior distribution like isotropic Gaussian noise. There exist several other generative models for 3D molecular structures, but none of them are designed for denoising. Typically, they generate equilibrium structures from scratch by iteratively adding new atoms [81, 82, 83, 84, 85] or by transforming samples from a prior distribution to a target distribution in one shot [86, 87, 88, 89]. Furthermore, generative models have been used to sample conformations given molecular graphs as input [90, 91, 92, 93, 94, 95, 96]. A common drawback of all generative methods is that, unlike relaxation-based methods, equilibrium structures are generated from scratch, which makes it difficult to steer the generation towards desired structures.
@@
We test the performance of our method on QM7-X [97], a dataset containing 42 000 equilibrium structures found with third-order self-consistent charge density functional tight binding [98] (DFTB3) [99, 100, 101] and many-body dispersion (MBD) [102, 103] corrections. They cover all molecular graphs in GDB13 [104] with up to 7 heavy atoms and include the elements H, C, N, O, S, and Cl. For each equilibrium structure, 100 non-equilibrium structures generated via normal-mode displacements of the equilibrium geometry are reported, including DFT calculations for energies and forces at the PBE0+MBD level[102, 105, 106] with FHI-aims[107, 108]. In our experiments, we employ several baselines for comparison, including the classical force field MMFF94, the semiempirical method GFN2-xTB, as well as a MLFF model with an equivalent neural network backbone architecture, and show that MoreRed performs favourably. Specifically, MoreRed accurately maps non-equilibrium structures back to the data manifold of equilibrium structures that it has been trained on. This is despite being trained on two orders of magnitude fewer structures than the MLFF. Additionally, while MLFFs can only relax structures that are covered by the distribution of training data, the inherent augmentation of the training data through the diffusion process enhances the robustness of MoreRed against variations in the noise distribution of the non-equilibrium test structures. Consequently, MoreRed successfully identifies the correct equilibrium structures for non-equilibrium inputs where MLFFs fail. We also show that the difference in DFT energies between the reference equilibrium structures in QM7-X and the structures obtained by MoreRed through molecular relaxation falls below the threshold of chemical accuracy.
@@ 2.1 Diffusion models
for $t=T,T-1,\dots,1$, which progressively removes the noise from the sample to denoise it, in the optimal case ending with a sample $\mathbf{x}_{0}$ from the target $q_{\mathrm{data}}(\mathbf{x}_{0})$ after $T$ reverse steps (see Figure 2, learned reverse generative process, right to left). This mimics molecular structure optimization by following atomistic forces. Here, the noise prediction $\bm{\varepsilon}_{\theta}(\mathbf{x}_{t},t)$ defines the opposite of the force directions that minimize the energy and the scaling terms that depend on $\alpha_{t}$ and $\beta_{t}$ determine the magnitudes and the step sizes used at each optimization step. The noise $\bar{\bm{\varepsilon}}$ added at each denoising step results in a stochastic optimizer instead of a deterministic one, which can be helpful in the case of the presence of many shallow local minima. However, because noise prediction $\bm{\varepsilon}_{\theta}(\mathbf{x}_{t},t)$ requires the diffusion time step $t$ as an input, diffusion models can only be used on structures with known noise level. For data generation, they start with samples from the known noise distribution at $t=T$, i.e., input structures that are pure noise. For molecular relaxation, in contrast, the non-equilibrium input structure can have an arbitrary level of perturbation such that the suitable initial time step is unknown, making the application of the standard plain diffusion models infeasible.
@@ 2.2 MoreRed: molecular relaxation by reverse diffusion
We design our diffusion model such that the distribution is invariant with respect to rotations $\mathcal{R}(\cdot)$ and translations $\mathcal{T}(\cdot)$, i.e., $p_{\theta}(\mathcal{T}(\mathcal{R}(\mathbf{x}_{t}))=p_{\theta}(\mathbf{x}_{t})$. To guarantee translational invariance, we center the atomic positions after each forward or reverse step. As proven by Xu et al [78], rotational invariance of the marginal distributions $p_{\theta}(\mathbf{x}_{t})$ is achieved by using an invariant prior $p_{T}(\mathbf{x}_{T})=\mathcal{N}(\bm{0},\mathbf{I})$, and an equivariant transition probability $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, which amounts to using an equivariant noise model $\bm{\varepsilon}_{\theta}(\mathbf{x}_{t},t)$. Therefore, we adopt the equivariant message passing architecture PaiNN [112], which allows us to directly predict equivariant tensor properties, such as the noise $\bm{\varepsilon}_{\theta}(\mathbf{x}_{t},t)$, as well as invariant scalar properties.
@@ 2.3 Diffusion time step prediction
where $\Gamma(\cdot)$ denotes the Gamma function. As discussed in Bishop et al [113], for large $D$, $p(\tilde{r})$ has a sharp peak at $\hat{r}\approx 1$, as illustrated in the left plot in Figure 7 in the supplementary information. This implies that each Gaussian component of $q_{t}$ at time step $t$ represents a sphere centered at a training sample $\mathbf{x}_{0}^{(i)}$, i.e. its density, represented by the set of diffused samples $\mathbf{x}_{t}^{(i)}$, is concentrated in a thin shell at radius $\hat{r}\approx 1$. Therefore, most of the samples $\mathbf{x}_{t}^{(i)}$ have similar distance from $\mathbf{x}_{0}^{(i)}$. Accordingly, assuming that the model can learn the data manifold, the distance, which corresponds to the noise level $(1-\bar{\alpha}_{t})$, is easy to identify from a single sample $\mathbf{x}_{t}^{(i)}$, as long as the noise level is small such that the mixture components (spheres) do not overlap. When the noise level increases, the diffused samples $\mathbf{x}_{t}^{(i)}$ from different training data points $\mathbf{x}_{0}^{(i)}$ overlap with each other. This overlap makes the estimation of the diffusion time step difficult, as indicated by the right plot in Figure 7. Further explanation and discussion are provided in section A.1 in the supplementary information, where we also show empirical evidence together with the derivation of Eq. (7).
@@ 2.4 Variants of reverse diffusion
A third variant, MoreRed joint training (MoreRed-JT), uses the same adaptive reverse diffusion process as described in Eq. (8) for MoreRed-AS but differs in the model definition and training. For MoreRed-AS we employ two separate neural networks with separate backbone representations, where one is used to predict the noise $\bm{\varepsilon}_{\theta}$ and the other one to predict the time step $\tau_{\Theta}$. For MoreRed-JT, we use one neural network as a shared backbone representation, and we add two prediction heads on top, one for the noise and one for the time step. This forces the noise and the time step heads to learn a joint molecular representation. We train this joint network by minimizing the joint loss $L_{\mathrm{joint}}=\eta L_{\mathrm{DDPM}}+(1-\eta)L_{\mathrm{DTP}}$, for $\eta\in[0,1]$ defining a trade-off between the two losses and combining Eq. 4 and Eq. 6. In the supplementary information, we provide details for the training in Algorithm (1) and for the sampling in Algorithm (2) in section A, and for the models in section C.1.
@@ 3.1 Diffusion time step prediction performance
Figure 3: a: Scatter plots of the RMSD of 10 000 non-equilibrium test structures from QM7-X and their equilibrium structures vs. the initial diffusion time step $\hat{t}$ predicted by MoreRed. Top: MoreRed-JT, where the time step and the noise are jointly predicted by the same neural network. Bottom: MoreRed-ITP/-AS, where the time step is predicted separately from the noise model with an individual neural network. b: Comparison of time step trajectories for relaxation of three non-equilibrium test structures from QM7-X in purple, pink, and orange. MoreRed-JT follows an adaptive schedule (the time step and noise are predicted at each denoising step; solid lines) and MoreRed-ITP follows a fixed schedule (the time step is only predicted at the initial denoising step; dashed lines). The dots show the predicted (but not utilized) time steps for MoreRed-ITP during denoising with the fixed schedule. The inset box shows the RMSDs between the reference equilibrium geometry and the non-equilibrium structure after and before relaxation, respectively.
@@
The diffusion time step predictor determines the starting step $\hat{t}$ of the reverse diffusion for molecular relaxation. The further a non-equilibrium structure deviates from its equilibrium structure, the more denoising steps are required, which means that a higher starting step $\hat{t}$ should be predicted. As can be seen in Figure 3a, the predicted starting steps correlate well with the RMSDs between the non-equilibrium and equilibrium test structures from QM7-X. This is a notable insight, as the non-equilibrium structures in QM7-X stem from DFTB normal-mode displacements of equilibrium geometries. All non-equilibrium examples used for training the time predictors, on the other hand, stem from diffusing equilibrium structures with Gaussian noise. Nevertheless, the time step predictors reliably predict $\hat{t}>0$ for all the 10 000 test structures, highlighting the robustness in identifying non-equilibrium structures even if they do not contain Gaussian noise. Moreover, we observe that a joint model for predicting both time step and noise, as in MoreRed-JT, leads to fewer outliers in the predictions of $\hat{t}$ and, consequently, a higher Pearson correlation coefficient ($\rho=0.85$; top) than the separately trained time step predictor in MoreRed-ITP/AS ($\rho=0.73$; bottom).
@@
Usually, diffusion models follow a fixed schedule where the time step $t$ is reduced by one after each denoising step until it reaches $t=0$. In MoreRed-ITP, we follow such a fixed schedule and only use the time step predictor once to obtain a suitable starting step $\hat{t}$. As described in Eq. (8) before, the other two variants, MoreRed-AS/-JT, utilize the time step predictor at every denoising step to obtain a new time step estimate $\hat{t}$. This results in an adaptive schedule, where the relaxation ends after a variable number of steps, as soon as $\hat{t}=0$ is predicted. Figure 3b shows the merit of this approach, where the time step trajectories during geometry relaxations of three different test structures are plotted in purple, pink, and orange. When following the fixed schedule (dashed lines), errors can occur and accumulate [111]. If not corrected, they lead to a mismatch between the true noise level in the structure and the time step $t$. Therefore, the relaxation may end before the sample reaches the equilibrium geometry. This can be observed for all three examples in the plot: The predicted but not enacted time step values associated with the fixed schedule (dotted lines) show a high value when the denoising with the fixed schedule ends, as the dashed line reaches $t=0$. In contrast, the adaptive schedule (solid lines), MoreRed-JT in this case, can account for such errors by adapting $\hat{t}$ at each denoising step. After converging to $\hat{t}=0$, the relaxed structures are significantly closer to the ground truth equilibrium geometry than the ones obtained with the fixed schedule (see the RMSD values after relaxation in the right-hand side box of Figure 3b).
@@
To conclude, we find that the time step predictor accurately identifies non-equilibrium structures, where larger time steps are predicted if the RMSD from the equilibrium structure is larger. Moreover, employing an adaptive schedule that utilizes the time step predictor to determine the time step at every denoising iteration proves beneficial compared to constantly decreasing the time step at a fixed rate. We provide further experiments in section B.5 in the supplementary information, where we show that the time step predictor also significantly enhances the performance of diffusion models in the original task of novel structure generation from complete noise.

### v2

@@ Abstract
Molecular relaxation, finding the equilibrium state of a non-equilibrium structure, is an essential component of computational chemistry to understand reactivity. Classical force field (FF) methods often rely on insufficient local energy minimization, while neural network FF models require large labeled datasets encompassing both equilibrium and non-equilibrium structures. As a remedy, we propose MoreRed, molecular relaxation by reverse diffusion, a conceptually novel and purely statistical approach where non-equilibrium structures are treated as noisy instances of their corresponding equilibrium states. To enable the denoising of arbitrarily noisy inputs via a generative diffusion model, we further introduce a novel diffusion time step predictor. Notably, MoreRed learns a simpler pseudo potential energy surface (PES) instead of the complex physical PES. It is trained on a significantly smaller, and thus computationally cheaper, dataset consisting of solely unlabeled equilibrium structures, avoiding the computation of non-equilibrium structures altogether. We compare MoreRed to classical FFs, equivariant neural network FFs trained on a large dataset of equilibrium and non-equilibrium data, as well as a semi-empirical tight-binding model. To assess this quantitatively, we evaluate the root-mean-square deviation between the found equilibrium structures and the reference equilibrium structures as well as their energies.
@@ 1 Introduction
Because iterative ab initio electronic structure calculations are computationally expensive and not feasible in high-throughput settings, methods for finding equilibrium structures need to be efficient [41, 42]. To address this limitation, numerous approaches have been developed that speed up the computation of forces but also suffer from a loss in accuracy compared to ab initio methods. These include classical force field (FF) methods like MMFF94 [43], the universal FF [44], or CHARMM [45], on the one hand, and semiempirical methods such as GFN2-xTB [46], PM6-7 [47, 48], or OM2 [49] on the other hand. Furthermore, machine learning FF (MLFF) models [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65] have emerged as promising alternatives to physical models. When trained on a sufficient amount of ab initio calculations, MLFF models, using kernel methods such as sGDML [66, 67, 68] or neural networks such as SchNet [69, 70], learn the physical PES and very efficiently predict the forces. Moreover, they have been shown to produce promising results in relaxation tasks [71, 72], while reducing computational cost by several magnitudes. Although MLFF models significantly accelerate gradient computations for structural relaxation, the training dataset must cover a wide range of the chemical space including equilibrium and non-equilibrium structures with accurately computed physical labels, introducing a large computational cost for generating a training dataset.
@@
An emerging machine learning-based approach to exploring chemical space is training generative models on a dataset of equilibrium structures in order to learn to generate new molecular structures. For instance, diffusion models have been used recently in molecule generation [73, 74, 75, 76, 77], conformer search [78] and molecular graph generation [79, 80]. They generate samples via iterative denoising, starting from a simple prior distribution like isotropic Gaussian noise. Several other generative models exist for 3D molecular structures, but they are generally not designed for denoising or generation from arbitrary states. Typically, they generate equilibrium structures from scratch by iteratively adding new atoms [81, 82, 83, 84, 85] or by transforming samples from a prior distribution to a target distribution in one shot [86, 87, 88, 89]. Furthermore, generative models have been used to sample conformations given molecular graphs as input [90, 91, 92, 93, 94, 95, 96]. A common drawback of conventional generative methods is that, unlike relaxation-based methods, equilibrium structures are generated from scratch, which makes it difficult to steer the generation towards desired structures. We define denoising as generating samples from the data manifold given arbitrarily noisy inputs. This task is fundamentally different from previous work [97, 98, 99, 100, 101] based on the idea of denoising autoencoders [102], where different denoising, yet not generative, techniques are used on molecular structures as an auxiliary or pre-training task for the original regression task of predicting forces and other molecular properties, aiming to improve data efficiency, generalizability and robustness. Relative to our approach, Hsu et al [103] use a score-based model to eliminate thermal noise or perturbations in the atomic positions of condensed materials, dealing with relatively small noise magnitudes.
@@
We test the performance of our method on QM7-X [104], a dataset containing 42 000 equilibrium structures found with third-order self-consistent charge density functional tight binding [105] (DFTB3) [106, 107, 108] and many-body dispersion (MBD) [109, 110] corrections. They cover all molecular graphs in GDB13 [111] with up to 7 heavy atoms and include the elements H, C, N, O, S, and Cl. For each equilibrium structure, 100 non-equilibrium structures generated via normal-mode displacements of the equilibrium geometry are reported, including DFT calculations for energies and forces at the PBE0+MBD level[109, 112, 113] with FHI-aims[114, 115]. In our experiments, we employ several baselines for comparison, including the classical FF MMFF94, the semiempirical method GFN2-xTB, as well as a MLFF model with an equivalent neural network backbone architecture, and show that MoreRed performs favourably. Specifically, MoreRed accurately maps non-equilibrium structures back to the data manifold of equilibrium structures that it has been trained on. This is despite being trained on two orders of magnitude fewer structures than the MLFF. Additionally, while MLFFs can only relax structures that are covered by the distribution of training data, the inherent augmentation of the training data through the diffusion process enhances the robustness of MoreRed against variations in the noise distribution of the non-equilibrium test structures. Consequently, MoreRed successfully identifies the correct equilibrium structures for non-equilibrium inputs where MLFFs fail. We also show that the difference in DFT energies between the reference equilibrium structures in QM7-X and the structures obtained by MoreRed through molecular relaxation falls below the threshold of chemical accuracy.
@@ 2.1 Diffusion models
for $t=T,T-1,\dots,1$, which progressively removes the noise from the sample to denoise it, in the optimal case ending with a sample $\mathbf{x}_{0}$ from the target $q_{\mathrm{data}}(\mathbf{x}_{0})$ after $T$ reverse steps (see Figure 2, learned reverse generative process, right to left). This mimics molecular structure optimization by following atomistic forces. Here, the noise prediction $\bm{\varepsilon}_{\theta}(\mathbf{x}_{t},t)$ defines the opposite of the force directions that minimize the energy and the scaling terms that depend on $\alpha_{t}$ and $\beta_{t}$ determine the magnitudes and the step sizes used at each optimization step. The noise $\bar{\bm{\varepsilon}}$ added at each denoising step results in a stochastic optimizer instead of a deterministic one, which can be helpful in the case of the presence of many shallow local minima. However, because noise prediction $\bm{\varepsilon}_{\theta}(\mathbf{x}_{t},t)$ requires the diffusion time step $t$ as an input, diffusion models can only be used on structures with known noise level. For data generation, they start with samples from the known noise distribution at $t=T$, i.e. input structures that are pure noise. For molecular relaxation, in contrast, the non-equilibrium input structure can have an arbitrary level of perturbation such that the suitable initial time step is unknown, making the application of the standard plain diffusion models infeasible.
@@ 2.2 MoreRed: Molecular relaxation by reverse diffusion
We design our diffusion model such that the distribution is invariant with respect to rotations $\mathcal{R}(\cdot)$ and translations $\mathcal{T}(\cdot)$, i.e. $p_{\theta}(\mathcal{T}(\mathcal{R}(\mathbf{x}_{t}))=p_{\theta}(\mathbf{x}_{t})$. To guarantee translational invariance, we center the atomic positions after each forward or reverse step. As proven by Xu et al [78], rotational invariance of the marginal distributions $p_{\theta}(\mathbf{x}_{t})$ is achieved by using an invariant prior $p_{T}(\mathbf{x}_{T})=\mathcal{N}(\bm{0},\mathbf{I})$, and an equivariant transition probability $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$, which amounts to using an equivariant noise model $\bm{\varepsilon}_{\theta}(\mathbf{x}_{t},t)$. Therefore, we adopt the equivariant message passing architecture PaiNN [119], which allows us to directly predict equivariant tensor properties, such as the noise $\bm{\varepsilon}_{\theta}(\mathbf{x}_{t},t)$, as well as invariant scalar properties.
@@ 2.3 Diffusion time step prediction
where $\Gamma(\cdot)$ denotes the Gamma function. As discussed in Bishop et al [120], for large $D$, $p(\tilde{r})$ has a sharp peak at $\hat{r}\approx 1$, as illustrated in the left plot in Figure 7 in the supplementary information. This implies that each Gaussian component of $q_{t}$ at time step $t$ represents a sphere centered at a training sample $\mathbf{x}_{0}^{(i)}$, i.e. its density, represented by the set of diffused samples $\mathbf{x}_{t}^{(i)}$, is concentrated in a thin shell at radius $\hat{r}\approx 1$. Therefore, most of the samples $\mathbf{x}_{t}^{(i)}$ have similar distance from $\mathbf{x}_{0}^{(i)}$. Accordingly, assuming that the model can learn the data manifold, the distance, which corresponds to the noise level $(1-\bar{\alpha}_{t})$, is easy to identify from a single sample $\mathbf{x}_{t}^{(i)}$, as long as the noise level is small such that the mixture components (spheres) do not overlap. When the noise level increases, the diffused samples $\mathbf{x}_{t}^{(i)}$ from different training data points $\mathbf{x}_{0}^{(i)}$ overlap with each other. This overlap makes the estimation of the diffusion time step difficult, as indicated by the right plot in Figure 7. Further explanation and discussion are provided in A.1 in the supplementary information, where we also show empirical evidence together with the derivation of Eq. (7).
@@ 2.4 Variants of reverse diffusion
A third variant, MoreRed joint training (MoreRed-JT), uses the same adaptive reverse diffusion process as described in Eq. (8) for MoreRed-AS but differs in the model definition and training. For MoreRed-AS we employ two separate neural networks with separate backbone representations, where one is used to predict the noise $\bm{\varepsilon}_{\theta}$ and the other one to predict the time step $\tau_{\Theta}$. For MoreRed-JT, we use one neural network as a shared backbone representation, and we add two prediction heads on top, one for the noise and one for the time step. This forces the noise and the time step heads to learn a joint molecular representation. We train this joint network by minimizing the joint loss $L_{\mathrm{joint}}=\eta L_{\mathrm{DDPM}}+(1-\eta)L_{\mathrm{DTP}}$, for $\eta\in[0,1]$ defining a trade-off between the two losses and combining Eq. 4 and Eq. 6. In the supplementary information, we provide details for the training in Algorithm (1) and for the sampling in Algorithm (2) in A, and for the models in C.1.
@@ 3.1 Diffusion time step prediction performance
Figure 3: a: Density plots of the RMSD of 10 000 non-equilibrium test structures from QM7-X and their equilibrium structures vs. the initial diffusion time step, $\hat{t}$, predicted by MoreRed-JT (top) and MoreRed-ITP/-AS (bottom). A brighter color (yellow) indicates a higher density of scatter points, a darker color (purple) indicates a lower density. For both, we show the Pearson correlation coefficient, $\rho$. The three crosses (red, orange, green) mark the structures for which a time step trajectory is shown in plot b in matching colors. b: Comparison of time step trajectories for relaxation of three non-equilibrium test structures from QM7-X in red, orange, and green (matching the crosses in plot a). MoreRed-JT follows an adaptive schedule (the time step and noise are predicted at each denoising step; solid lines) and MoreRed-ITP follows a fixed schedule (the time step is only predicted at the initial denoising step; dashed lines). The dots show the predicted (but not utilized) time steps for MoreRed-ITP during denoising with the fixed schedule. All two methods start from a comparable initial time step for each non-equilibrium structure. The inset box shows the RMSDs between the reference equilibrium geometry and the non-equilibrium structure after and before relaxation, respectively.
@@
The diffusion time step predictor determines the starting step $\hat{t}$ of the reverse diffusion for molecular relaxation. The further a non-equilibrium structure deviates from its equilibrium structure, the more denoising steps are required, which means that a higher starting step $\hat{t}$ should be predicted. As can be seen in Figure 3a, the predicted starting steps correlate well with the RMSDs between the non-equilibrium and equilibrium test structures from QM7-X. This is a notable insight, as the non-equilibrium structures in QM7-X stem from DFTB normal-mode displacements of equilibrium geometries. All non-equilibrium examples used for training the time predictors, on the other hand, stem from diffusing equilibrium structures with Gaussian noise. Nevertheless, the time step predictors reliably predict $\hat{t}>0$ for all the 10 000 test structures, highlighting the robustness in identifying non-equilibrium structures even if they do not contain Gaussian noise. Moreover, we observe that a joint model for predicting both time step and noise, as in MoreRed-JT, leads to fewer outliers in the predictions of $\hat{t}$ and, consequently, a higher Pearson correlation coefficient ($\rho=0.85$; top) than the separately trained time step predictor in MoreRed-ITP/AS ($\rho=0.72$; bottom).
@@
Usually, diffusion models follow a fixed schedule where the time step $t$ is reduced by one after each denoising step until it reaches $t=0$. In MoreRed-ITP, we follow such a fixed schedule and only use the time step predictor once to obtain a suitable starting step $\hat{t}$. As described in Eq. (8) before, the other two variants, MoreRed-AS/-JT, utilize the time step predictor at every denoising step to obtain a new time step estimate $\hat{t}$. This results in an adaptive schedule, where the relaxation ends after a variable number of steps, as soon as $\hat{t}=0$ is predicted. Figure 3b shows the merit of this approach, where the time step trajectories during geometry relaxations of three different test structures are plotted in red, orange, and green. When following the fixed schedule (dashed lines), errors can occur and accumulate [118]. If not corrected, they lead to a mismatch between the true noise level in the structure and the time step $t$. Therefore, the relaxation may end before the sample reaches the equilibrium geometry. This can be observed for all three examples in the plot: The predicted but not enacted time step values associated with the fixed schedule (dotted lines) show a high value when the denoising with the fixed schedule ends, as the dashed line reaches $t=0$. In contrast, the adaptive schedule (solid lines), MoreRed-JT in this case, can account for such errors by adapting $\hat{t}$ at each denoising step. After converging to $\hat{t}=0$, the relaxed structures are significantly closer to the ground truth equilibrium geometry than those obtained with the fixed schedule, despite starting from comparable initial time steps (see the RMSD values after relaxation in the right-hand side box of Figure 3b). In Figure 13 in the B.6 we present similar trajectories to Figure 3b but for three cases where the predicted initial time steps are very different for both methods. It highlights that even when MoreRed-ITP starts from a higher initial step, it is still outperformed by MoreRed-JT.
@@
To conclude, we find that the time step predictor accurately identifies non-equilibrium structures, where larger time steps are predicted if the RMSD from the equilibrium structure is larger. Moreover, employing an adaptive schedule that utilizes the time step predictor to determine the time step at every denoising iteration proves beneficial compared to constantly decreasing the time step at a fixed rate. We provide further experiments in B.6 in the supplementary information, where we show that the time step predictor also significantly enhances the performance of diffusion models in the original task of novel structure generation from complete noise.